OpenStudy (anonymous):

Find the area of the region in the first quadrant (upper right quadrant) bounded by the curves y = sin x cos2x, y = 2x cox(x2), and y = 4 – 4x.

OpenStudy (anonymous):

@mathmale hi! lol

OpenStudy (anonymous):

i just need help solving the b-a part

OpenStudy (anonymous):

i dont know which curves to use to find the boundaries

OpenStudy (mathmale):

"LOL" is right. This is a very challenging problem. I do see that x begins with 0 and that y also begins with 0 (due to the phrase "first quadrant."

OpenStudy (mathmale):

Secondly, I need to verify that I understand your presentation of these functions. Are the following correct or not?

OpenStudy (anonymous):

yes.

OpenStudy (mathmale):

\[ y = \sin x *\cos2x, y = 2*x *\cos(x^2)\]

OpenStudy (anonymous):

i got it! lower bound=0 upper bound is pi/4.

OpenStudy (mathmale):

I'm curious how you arrived at that upper bound. Was it given? I haven't actually done this problem, so don't know yet what my own answer would look like.